# Avenida Complutense 40, 28040. Madrid, Spain

### v1

@@ Abstract
We present a reinterpretation of selected ATLAS and CMS searches for Higgs boson decays into long-lived particles (LLPs) within the scalar portal framework, where the standard model is extended by a light scalar that mixes with the Higgs boson. Using public HEPData results, we evaluate the sensitivity of several experimental strategies targeting different final states. The analysis shows how published searches probe complementary regions of the model parameter space, and it provides a comparative assessment of their coverage. The study also identifies the existing gaps in phase space coverage and highlights key areas for future improvement in LLP searches, which will be crucial to maximize the discovery potential for exotic decays of the Higgs boson into LLPs. Finally, we extrapolate the constraints to the integrated luminosity of the High-Luminosity LHC under different assumptions and compare them with the projected sensitivities of proposed dedicated LLP experiments at CERN.
@@ 2 Model description: scalar portal
The non-zero $\mu$ leads to the mixing of $h$ and $S$ states, where $h$ is the physical $125~\text{GeV}$ Higgs boson. In the limit of small mixing it can be written as
@@ 4.1 LHC constraints
While this manuscript was in preparation, CMS updated the displaced $\mu^{+}\mu^{-}$ constraints for $m(\text{S})<5~\text{GeV}$ using the first 64.2 fb-1 of Run 3 scouting data CMS-PAS-EXO-24-016 , and released a new search targeting $\text{h}\to\text{S}\text{S}$ with mixed $\mu^{+}\mu^{-}$ and hadronic decays ($\pi^{\pm}\pi^{\mp}$ or $K^{\pm}K^{\mp}$CMS-PAS-EXO-24-034 , reaching ${\cal B}({\text{h}\to\text{S}\text{S}})\approx 10^{-4}$ at $m(\text{S})=0.6~\text{GeV}$. ATLAS also updated its displaced vertex search in the muon detectors ATLAS:2025pak . The addition of these new results is not expected to significantly change any of the conclusions of this paper.
@@ 4.2 HL-LHC projections
  • 3 ab-1 scaled by $\sqrt{\cal{L}}$ (green): current 95% CL observed upper limits are scaled by $\sqrt{140/3000}$. This corresponds to a pessimistic scenario in which backgrounds remain and no improvements in signal efficiency are achieved at the HL-LHC.
@@
In each HL-LHC projection, all upper limits are assumed to scale equally. In practice, however, some search strategies offer greater potential for R&D than others, which should be kept in mind when interpreting these projections. Because of the limited granularity of interpretations at the smallest values of $\sin\theta$ (e.g see Fig. 4 for $m(\text{S})<10~\text{GeV}$), additional points were obtained by fitting the constraints with a linear regression in log-log plane to produce smoother contours in Fig. 7. These HL-LHC projections are compared with the expected sensitivity estimates from Ref. PBC:2025sny for dedicated LLP detectors at CERN. Following the format in Ref. PBC:2025sny , the comparison is performed in the $\{m(\text{S}),\sin\theta\}$ plane instead of $\{m(\text{S}),\sin^{2}\theta\}$.
@@ 5 Conclusions
[…]

### v2

@@ Abstract
We present a reinterpretation of selected ATLAS and CMS searches for Higgs boson decays into long-lived particles (LLPs) within the scalar portal framework, where the standard model is extended by a light scalar that mixes with the Higgs boson. Using public HEPData results, we evaluate the sensitivity of several experimental strategies targeting different final states. The analysis shows how published searches probe complementary regions of the model parameter space, and it provides a comparative assessment of their coverage. The study also identifies the existing gaps in phase space coverage and highlights key areas for future improvement in LLP searches, which will be crucial to maximize the discovery potential for exotic decays of the Higgs boson into LLPs. Finally, we extrapolate the constraints to the integrated luminosity of the High-Luminosity LHC under different assumptions and compare them with the projected sensitivities of proposed dedicated LLP experiments at CERN, showing that these experiments will further extend the experimental reach independently of the assumed extrapolation.
@@ 2 Model description: scalar portal
The non-zero $\mu$ leads to the mixing of $h$ and $S$ states, where $h$ is the physical $125~\text{GeV}$ Higgs boson. In the limit of small mixing angle, it can be written as
@@ 4.1 LHC constraints
While this manuscript was in preparation, CMS updated the displaced $\mu^{+}\mu^{-}$ constraints for $m(\text{S})<5~\text{GeV}$ using the first 64.2 fb-1 of Run 3 scouting data CMS-PAS-EXO-24-016, and released a new search targeting $\text{h}\to\text{S}\text{S}$ with mixed $\mu^{+}\mu^{-}$ and hadronic decays ($\pi^{\pm}\pi^{\mp}$ or $K^{\pm}K^{\mp}$CMS-PAS-EXO-24-034, reaching ${\cal B}({\text{h}\to\text{S}\text{S}})\approx 0.1\%$ at $m(\text{S})=0.6~\text{GeV}$ for $c\tau(\text{S})=5$ mm. ATLAS also updated its displaced vertex search in the muon detectors ATLAS:2025pak. The addition of these new results is not expected to significantly change any of the conclusions of this paper.
@@ 4.2 HL-LHC projections
  • 3 ab-1 scaled by $\sqrt{\cal{L}}$ (green): current 95% CL observed upper limits are scaled by $\sqrt{140/3000}$. This corresponds to a conservative scenario in which backgrounds remain and no improvements in signal efficiency are achieved at the HL-LHC.
@@
In each HL-LHC projection, all upper limits are assumed to scale equally. In practice, however, some search strategies offer greater potential for R&D than others, which should be kept in mind when interpreting these projections. Since LHC LLP summaries are available only for 95% CL observed limits, all extrapolations are based on observed limits. Because of the limited granularity of interpretations at the smallest values of $\sin\theta$ (e.g see Fig. 4 for $m(\text{S})<10~\text{GeV}$), additional points were obtained by fitting the constraints with a linear regression in log-log plane to produce smoother contours in Fig. 7. These HL-LHC projections are compared with the expected sensitivity estimates from Ref. PBC:2025sny for dedicated LLP detectors at CERN. Following the format in Ref. PBC:2025sny, the comparison is performed in the $\{m(\text{S}),\sin\theta\}$ plane instead of $\{m(\text{S}),\sin^{2}\theta\}$.
@@ 5 Conclusions
All figures presented in this paper, expressed as functions of both $\sin\theta$ and $\sin^{2}\theta$, are publicly available in the GitHub  repository LLPPlotsRepo.